OpenStudy (anonymous):

what is 5g+h=g ;solve for g

OpenStudy (anonymous):

Make your g by itself so substract g from both sides.

OpenStudy (anonymous):

Then, substract h from both sides.

OpenStudy (anonymous):

oh thnx.

OpenStudy (anonymous):

wait, i am confused now

OpenStudy (anonymous):

Confused how?

OpenStudy (anonymous):

in my problem, how do i subtract g from both sides????

OpenStudy (anonymous):

can u write it out for me please?

OpenStudy (anonymous):

Think of the equation like a balance, see-saw, whatever.... you can do what you want to one side as long as you do the same thing to the other side... if you don't, it will "tip" and not be an equation any longer. So, if you want, subtract g from both sides... on the left, it would be "5g+h-g" --> 4g+h and on the right, it would just be "g-g" ---> 0

OpenStudy (anonymous):

oh, so it is no possible answer?

OpenStudy (anonymous):

So, to "get rid of the g" on the right side of the original equation, subtract g, but you have to do it from both sides.

OpenStudy (anonymous):

No, you can get an answer for g, but it will have an "h" in it. 5g+h=g simplifies to 4g+h=0 simplifies to 4g = h simplifies to g=(1/4)h

OpenStudy (anonymous):

Each step moves/modifies/whatever something from both sides of the equation. First, subtract g. Then simplify. Then subtract h from both sides. Then multiply both sides by (1/4).

OpenStudy (anonymous):

So the answer for g is that g = (1/4)h or another way would be g = h/4

OpenStudy (anonymous):

wait could u draw the steps?

OpenStudy (anonymous):

you don't get a number for g, you get an expression using h. but that's ok!

OpenStudy (anonymous):

my teacher wld kill me!

OpenStudy (anonymous):

um, hope not!

OpenStudy (anonymous):

could u draw the steps tho

OpenStudy (anonymous):

What I do is write the equation on one line. Then write the next "version" of it on the next line with the one thing I changed. 5g+h=g 5g+h-g=g-g 4g+h=0 4g+h-h=0-h 4g=-h g=-h/4 (note, I messed up earlier and left off the minus sign)

OpenStudy (anonymous):

THANK YOU SO MUCH!!

OpenStudy (anonymous):

hope you understand it too :)